OpenStudy (anonymous):

Given the function f(x) = 3x +3, find the value of f(1/3). 3 1/3 4 1

OpenStudy (jackellyn):

Substitute 1/3 anytime you see an x 3(1/3) +3 =?

hero (hero):

f(x) = 3x + 3 f(1/3) = 3(1/3)+3 f(1/3) = 1 + 3 f(1/3) = 4